OpenStudy (moongazer):

is 50*2+25-10MOD4+10 the same as (50 2+25-10) mod (4+10) ????

OpenStudy (anonymous):

that notation is a little weird. why would someone write mod (4+10) and not mod 15?

OpenStudy (anonymous):

Usually when you are working with modular arithmetic, you have to mod the entire statement. It doesnt make any sense to only look at a portion of the statement.

OpenStudy (moongazer):

so before you do MOD you need first solve the other operations?

OpenStudy (moongazer):

why is it 20MOD110 is 20?

OpenStudy (anonymous):

you can go either way. Do all the operations first and then mod, or mod completely first, then do operations. just dont mix up the two. and yes, 20 mod 110 is 20.

OpenStudy (moongazer):

how?20 mod 110 is 20

OpenStudy (moongazer):

could you show the solution

OpenStudy (anonymous):

There are a couple of ways to see it. One way to think of x mod 110 is what is the remainder when i divide by 110? well 20 is already less than 110, so its remainder would be 20. Another number equivalent to 20 mod 110 is 130. Because when i divide 130 by 110, i get a remainder of 20.

OpenStudy (moongazer):

what? i already know that MOD is the remainder in division but i can't understand how you divided 20 to 100 sorry if i did'nt understood your explanation

OpenStudy (anonymous):

are we working with 100? or 110?

OpenStudy (moongazer):

110

OpenStudy (moongazer):

typo :)

OpenStudy (anonymous):

you would divide 20 by 110. The answer is 0 with a remainder of 20. Or you can write out 20 using the division algorithm: 20 = 0*110+20 The remainder is 20 so 20 mod 110 is 20. In fact any number that is less than 110 will be equivalent to itself. 35 mod 110 is 35. 109 mod 110 is 109. etc.

OpenStudy (moongazer):

could you show me on how you divide 20 by 110 and get 20?

OpenStudy (anonymous):

you are getting a remainder of 20. Not 20. Im not saying: \[\frac{20}{110} = 20\] thats wrong. Im saying: \[\frac{20}{110} = 0+\frac{20}{110}\] so the remainder is 20.

OpenStudy (anonymous):

if i did 130 it would be: \[\frac{130}{110} = 1+\frac{20}{110}\] so the remainder is 20. so 130 mod 110 = 20

OpenStudy (moongazer):

yup saying it to remainder

OpenStudy (moongazer):

how did you get remainder 20?

OpenStudy (moongazer):

sorry if i can't really understand

OpenStudy (anonymous):

Maybe its because we are working with big numbers. Could you tell me what the remainder is when i divide 15 by 6?

OpenStudy (moongazer):

is my insight correct that if the left part of the mod is less than the right part, the mod will always be the left part? e.g 30mod3425234534 mod will be 30?

OpenStudy (anonymous):

yes that is right. You need to be able to calculate remainders though. otherwise if the number is bigger than the mod you wont know how to work it out.

OpenStudy (moongazer):

because mod is always 30 in wolfram alpha if i enter a larger number on the right side of mod

OpenStudy (moongazer):

so 60mod26283475908273459082734509287345(or whatever number but larger than 60) the Mod will always be 60?

OpenStudy (anonymous):

yes

OpenStudy (moongazer):

thanks